### v2

Abstract
Active learning strategies aim to train high-performance models with minimal labeled data by selecting the most informative instances for labeling. However, existing methods for assessing data informativeness often fail to align directly with task model performance metrics, such as mean average precision (mAP) in object detection. This paper introduces Mean-AP Guided Reinforced Active Learning for Object Detection (MGRAL), a novel approach that leverages the concept of expected model output changes as informativeness for deep detection networks, directly optimizing the sampling strategy using mAP. MGRAL employs a reinforcement learning agent based on LSTM architecture to efficiently navigate the combinatorial challenge of batch sample selection and the non-differentiable nature between performance and selected batches. The agent optimizes selection using policy gradient with mAP improvement as the reward signal. To address the computational intensity of mAP estimation with unlabeled samples, we implement fast look-up tables, ensuring real-world feasibility. We evaluate MGRAL on PASCAL VOC and MS COCO benchmarks across various backbone architectures. Our approach demonstrates strong performance, establishing a new paradigm in reinforcement learning-based active learning for object detection.
Index Terms:
active learning, object detection, reinforcement learning, policy gradient.
I Introduction
The pursuit of artificial intelligence fundamentally revolves around optimizing two key elements: data and models. While significant advancements have been made in refining model architectures, the focus of contemporary research is increasingly shifting towards more efficient data utilization strategies. Among these, Active Learning (AL) stands out for its ability to efficiently train high-performance models with minimal labeled data. This approach is particularly valuable in environments where there is a continuous flow of operational data requiring high labeling costs. By strategically selecting and annotating the most informative ones, AL optimizes data utilization, significantly boosting the efficiency of AI systems.
Determining the most informative data typically involves identifying data points that are complementary to the currently labeled ones within the model’s feature space, primarily characterized by uncertainty. Although various definitions of uncertainty have been proposed and utilized as query strategies in active learning [1, 2, 3, 4, 5], these do not always correlate directly with the performance metrics of the task model, such as mean Average Precision (mAP) in object detection. EMOC [6] proposed the expected model output change to measure sample informativeness, but this approach was limited to Gaussian Process Regression models and lacks applicability to deep learning architectures. Most existing methods, including EMOC, evaluate samples individually, overlooking the collective impact of sample batches in AL scenarios.
Figure 1: Training and selection pipelines of Mean-AP Guided Reinforced Active Learning (MGRAL) for object detection. (a) Data sampling agent training phase. An RL-based agent is trained using $\Delta$mAP as reward to prioritize data batch that improves detection neural network performance most. (b) Data selection with trained sampling agent. The trained agent is used as the query strategy to select the most informative data batch for annotation and detection model fine-tuning.
In this paper, we introduce Mean-AP Guided Reinforced Active Learning (MGRAL), a novel approach that directly leverages mAP to guide data sampling for object detection. The core challenge lies in the discrete nature of batch selection from the unlabeled pool, making the relationship between selected samples and mAP changes non-differentiable. To address this, we employ a reinforcement learning agent that uses mAP variation ($\Delta$mAP) as the reward, optimizing the selection process through policy gradient techniques. This approach enables efficient exploration of possible batch combinations, maximizing mAP improvement per selected batch.
To estimate the potential change on mAP when including unlabeled samples, MGRAL employs an unsupervised model. This approach is effective as it captures global trends sufficient for identifying potentially impactful data. The unsupervised model’s output need not be absolutely correct labels; rather, it aims to correctly estimate which data significantly differ from the model’s current decision boundaries or distributions. This relative information is crucial for guiding the active learning process in selecting the most informative samples for labeling. Additionally, to address the extensive training time required by the reinforcement learning agent, which necessitates retraining the semi-supervised model iteratively, we implement fast lookup tables for acceleration. Empirical results demonstrate our method’s efficiency on both Pascal VOC [7] and MS COCO [8] datasets.
Our contributions can be summarized as: (1) We propose MAGRAL, a novel approach that extends the concept of expected model output changes to deep neural networks for object detection, directly aligning the sampling strategy with mAP. (2) We introduce a reinforcement learning agent for efficient batch sample selection, addressing the combinatorial explosion and non-differentiable challenge and optimizing for mAP improvement. (3) We implement practical techniques, including an unsupervised model for mAP estimation and a fast look-up table, making our method feasible for real-world active learning scenarios in object detection.
II Related Works
II-A Performance-Driven Active Learning Strategies
Classical active learning methods typically fall into uncertainty-based, distribution-based, and hybrid categories. Uncertainty-based methods (e.g., information theoretic heuristics [9], query-by-committee [10, 11], and Bayesian models [12, 13]) prioritize perplexing data points but may lead to redundancy. Distribution-based approaches like Core-set [14] aim for diversity but struggle with high-dimensional spaces [15, 16]. Hybrid methods [17, 9, 18, 19, 5] attempt to balance both, yet face challenges in effectively combining these metrics. However, these methods may not directly align with the task model’s performance.
Freytag et al. [6] proposed expected model output changes to measure sample informativeness, but it was limited to Gaussian process regression models and lacks applicability to deep learning architectures. Most previous methods focus on individual sample selection, neglecting batch selection importance. Our work extends these concepts to deep neural networks, directly aligning sampling with mAP for object detection and addressing batch selection challenges.
II-B Deep Learning and Reinforcement Learning in AL
Deep learning has introduced methods that calculate informativeness by learning or use neural network-driven selection, such as learning loss estimation [1] and adversarial-based VAAL [20]. Meanwhile, reinforcement learning has been adopted to learn better query strategies, using techniques like policy gradient methods [21], imitation learning [22], bi-directional RNNs [23], and Deep Q-Networks [24, 25]. However, these methods often struggle with batch sample selection and integrating multi-instance uncertainty within a single image for object detection tasks. Our approach addresses these challenges by using a reinforcement learning agent optimized for batch selection with $\Delta\text{mAP}$ as the reward.
II-C Active Learning for Object Detection
Recent advances in active learning for object detection include LL4AL [1], which adapts instance loss predictions, and Aghdam et al. [26], which combines uncertainty metrics for foreground objects and background pixels. CDAL [2] enhances sample representativeness through spatial context, while MIAL [3, 27] employs adversarial classifiers and a semi-supervised framework. EBAL [5] integrates uncertainty and diversity but faces challenges with computational complexity and class imbalance. Our method differs by directly utilizing mAP to guide the selection process, addressing the limitations of previous approaches in balancing various metrics and handling batch selection efficiently.
III Methodology
III-A Problem Definition
Active learning for object detection follows the setting that a small labeled set $X_{L}$ containing images with instance labels, denoted as $\{(x_{L},y_{L})\}$ and a large unlabeled set $X_{U}$ without labels, denoted as $\{(x_{U})\}$ are given. The labels include locations of bounding boxes and their corresponding categories.
For each cycle of the active learning process, a detection model $M_{i}$ ($i$ denotes the cycle subscript) is initially trained using the labeled dataset $X_{L}^{i}$. Subsequently, active learning employs a query strategy to select a subset of images $X_{S}^{i}=\{(x_{S}^{i})\}$ from the unlabeled dataset $X_{U}^{i}$. These images are then annotated and integrated into $X_{L}^{i}$ to create an updated labeled dataset $X_{L}^{i+1}=X_{L}^{i}\cup\{(x_{S}^{i},y_{S}^{i})\}$. The updated dataset $X_{L}^{i+1}$ is used for training the next iteration of the detection model $M_{i+1}$. This cycle repeats until the size of the labeled dataset reaches the predefined budget $B$. The effectiveness of the query strategy is pivotal, as it directly influences the enhancement of model performance with each cycle, motivating the development of our proposed method.
III-B Overview of MGRAL Pipeline
Our proposed method MGRAL derives its query strategy by utilizing mean Average Precision of the task learner to maximize the impact of selected batches.
As shown in Figure 1, MGRAL integrates a reinforcement learning-based sampling agent (controller) into the typical pool-based active learning pipeline. This agent addresses a fundamental challenge in active learning for object detection: the difficulty in directly optimizing batch selection based on mAP improvement. In conventional approaches, it’s challenging to establish a direct link between the selected batch of samples and the resulting mAP improvement due to the discrete and combinatorial nature of batch selection. The reinforcement learning agent overcomes this by learning to navigate this complex, discrete search space, using mAP improvements as feedback. This approach allows the agent to effectively learn the correlation between batch selections and their impact on mAP, without requiring a differentiable relationship.
Figure 2: Data sampling agent architecture.
III-C MGRAL Data Sampling Agent
The MGRAL data sampling agent employs a Long Short-Term Memory (LSTM) [28] network to optimize the selection of informative images from the unlabeled pool. As shown in Figure 2, each image is first processed through a feature extraction network $\Phi$, which derives a feature vector using a pre-trained detection model.
The LSTM Modules operate in a parameter-sharing mode across all images, enhancing scalability and preventing gradient vanishing. Moreover, we compute the embedding vector for each image by concatenating the feature vector with the decision vector from the preceding unit. This approach integrates both new data and historical decisions. The interaction between the LSTM Module outputs and the decision-making process is depicted as:
[…]
$A_{i}$ $=\Psi_{i}(h_{i}),\vspace{-5pt}$ (2)
where []delimited-[][$\cdot$][ ⋅ ] denotes concatenation operation, $\Psi_{i}$ is the decoding network for the $i$-th image. This structure ensures each image’s selection considers its potential contribution to the model’s performance, as quantified by changes in mAP.
III-D Performance-Driven Reward Design
The core of our approach lies in directly associating data sampling decisions with the performance improvements of the detection model. We define the agent’s reward as the variation in mean Average Precision ($\Delta\text{mAP}$). For each training iteration $i$, we combine the current labeled dataset $X_{L}$ with the newly selected unlabeled samples $X_{S}^{i}$. A new detector is then trained on this combined set with $\text{mAP}_{i}$ computed. The reward is calculated as $\Delta\text{mAP}=\text{mAP}_{i}-\text{mAP}_{i-1}$.
Given that newly selected samples are initially unlabeled, we employ a semi-supervised detection model to estimate the detector’s performance. This proxy model facilitates the estimation of mAP in the absence of fully labeled data, crucial for our active learning setting. The semi-supervised model captures global trends of unlabeled data distribution sufficient for identifying potentially impactful data, providing relative information about data informativeness rather than accurate predictions or pseudo-labels. Importantly, even if the semi-supervised model’s performance is not optimal, it does not significantly impact our method’s effectiveness as we only require a reasonable ranking of data informativeness. While semi-supervised models cannot completely replace supervised learning, this active learning approach allows us to efficiently select valuable data for annotation, optimizing the trade-off between labeling effort and model performance.
III-E Training Process and Optimization
The training process of MGRAL involves iterative refinement of the data sampling agent. In each iteration, the agent first selects a candidate set from the unlabeled pool, which is then used to update the semi-supervised detection model. The model’s performance, measured by estimated mAP, guides the agent’s learning. To stabilize training, we employ a moving average mAP (mAPref) as a baseline:
$\text{mAP}_{ref}=\lambda*\text{mAP}_{ref}+(1-\lambda)*(\text{mAP}_{i}-\text{mAP}_{ref}).$ (3)
The agent’s loss is calculated as the negative difference between the current mAP and this reference:
$loss_{\text{agent}}=-(\text{mAP}_{i}-\text{mAP}_{ref}).$ (4)
This process repeats for $max\_search\_iters$ times, progressively improving the agent’s sample selection strategy.
III-F Acceleration Technique
MGRAL requires frequent retraining of the semi-supervised detector to estimate mAP for each potential batch selection, leading to significant computational costs. To mitigate this extensive training duration, we employ a fast lookup table technique. We pre-compute and store model performances for a series of labeled dataset sizes, incrementing by the batch size used in each active learning cycle, from the initial set up to budget $B$. During training, instead of retraining the semi-supervised model for each new batch selection, we estimate mAP by comparing the visual features of the selected batch to those in the lookup table.
Specifically, we use L2 distance (outperform cosine similarity) to measure similarity between the visual representations of the selected batch and the pre-computed datasets. The mAP is approximated through a weighted summation of the most similar pre-computed results, with weights inversely proportional to the measured distances to account for confidence. If all similar pre-computed results exceed a set distance threshold (the mean feature distance minus one standard deviation), we revert to training the semi-supervised model directly for an accurate mAP. This approach significantly accelerates the training process while maintaining accuracy when needed.
(a) SSD on Pascal VOC
(b) RetinaNet on MS COCO
Figure 3: Comparative performance of active learning methods.
TABLE I: Training time of MGRAL with and without lookup table on VOC.
Method Time of 10 Iters Time for One Cycle
w/o acceleration 4800 min unknown
w/ lookup table 3 min 640 min
TABLE II: Overall time cost comparison. First active learning cycle on VOC.
Method Training Time Inference Time
Entropy 0 5 min
MIAL [3] 7 h 13 min 43 min
EBAL [5] 0 181 min
Ours (9 h) + 7 h 8 min 0.5 min
IV Experiments
IV-A Experiment Details
Datasets: We evaluate MGRAL on PASCAL VOC [7] (VOC07+12 trainval: 16,551 images, 20 classes; VOC07 test: 4,952 images) and MS COCO [8] (train2017: 117k images; val2017: 5k images; 80 classes).
Active Learning Settings: For VOC, we follow [1, 3, 5] with an initial pool of 1,000 labeled images and select 1,000 images per cycle for 10 cycles. For COCO, we start with 2.0% labeled data, adding 2.0% per cycle until reaching 10.0%.
Detector and Semi-supervise Settings: On PASCAL VOC, we use SSD [29] detector with ISD-SSD [30], training for 300 epochs (learning rate: 0.001, reduced to 0.0001 after 240 epochs; batch size: 32). For COCO, we employ RetinaNet [31] detector with SED-SSOD [32] for 180k iterations (batch size: 8; learning rate: 0.02, dropped by 0.1 at the 120k and 160k).
Data Sampling Agent Configuration: The controller uses a 257-dimensional vector (256 for image embeddings, 1 for previous selection score) with an LSTM of the same size. We optimize using Adam [33] (learning rate: $3.5\times 10^{-4}$).
RL Training Settings: The maximum search iterations for each cycle are set to 2,200 for Pascal VOC and 800 for MS COCO, with a baseline decay factor of 0.5 for mAPref.
Fast Lookup Table: For Pascal VOC, we pre-compute 200 experiments per cycle using ISD-SSD. For COCO, we use 30 experiments per cycle with SED-SSOD, totaling 150 records.
IV-B Overall Performance
We compare MGRAL with various baselines including random sampling, entropy sampling, Core-set [14], CDAL [2], LL4AL [1], MIAL [3, 27], and the state-of-the-art EBAL [5]. Experiments were conducted on GTX 1080Ti GPUs for PASCAL VOC and Tesla V100 GPUs for MS COCO.
As shown in Fig.3, MGRAL consistently outperforms all methods on PASCAL VOC, demonstrating the effectiveness of our mAP-guided approach. On MS COCO, MGRAL beats all other baselines across all cycles except EBAL [5]. MGRAL exhibits the steepest performance curve, ultimately surpassing EBAL in the final cycle. This curve indicates the mAP-guided strategy and direct mAP optimization by policy gradient offer unique advantages in active learning for object detection.
Figure 4: Selected samples of first cycle on Pascal VOC.
IV-C Qualitative Analysis
Fig.4 compares first-cycle selections across query strategies. Mean entropy sampling (top left) favors exposed or blurred images, potentially hindering initial training. Sum entropy sampling (top right) selects images with multiple instances, but often from redundant categories. MIAL[3] (bottom left) and MGRAL (bottom right) prefer single, centrally-located objects, providing clearer learning signals. Notably, MGRAL shows a stronger preference for diverse categories, likely enhancing detector robustness across classes. This diversity-oriented selection explains MGRAL’s superior mAP performance, demonstrating the effectiveness of its mAP-guided approach in creating a balanced and effective initial dataset.
IV-D Efficiency and Ablation Analysis
The fast lookup table significantly reduces training time from 4800 to 3 minutes for ten iterations on four GTX 1080Ti GPUs (Tab.3). Compared to entropy sampling, MIAL [3], and EBAL [5], MGRAL shows competitive overall time efficiency on VOC, offering shorter inference times in subsequent cycles despite a longer initial setup (Tab.3). These demonstrate MGRAL’s long-term efficiency in active learning, effectively balancing setup costs with improved subsequent performance.
V Conclusion
We presented MGRAL, a Mean-AP Guided Reinforced Active Learning for object detection. By using $\Delta\text{mAP}$ as reward, our method aligns batch selection with performance improvement, addressing the non-differentiable nature of this process. MGRAL demonstrates strong results on VOC and COCO. While our fast lookup table effectively accelerates training, it remains a preliminary solution open for further optimization. This work establishes a new paradigm in active learning for object detection, demonstrating the effectiveness of performance-driven reinforcement learning in this area.